OpenStudy (megannicole51):

A rod with density 3+sin(x) lies on the x-axis between x=0 and x=pi. Find the mass and center of mass of the rod.

OpenStudy (anonymous):

so here you integrate... from 0 to pi. \[m = \int\limits_{0}^{\pi}(3+\sin x)dx\] and the center of mass of the rod is \[CM = \frac{ 1 }{ m }\int\limits_{0}^{\pi}x(3+\sin x)dx\]

OpenStudy (anonymous):

you there?

OpenStudy (megannicole51):

yeah im trying to figure it out sorry

OpenStudy (anonymous):

no it's okay. take your time

OpenStudy (megannicole51):

yay i got the right answers! :D thank you!

OpenStudy (anonymous):

you're welcome!

OpenStudy (megannicole51):

how do you find a mass of a triangle? is it the same or different?

OpenStudy (anonymous):

what's the density function?

OpenStudy (megannicole51):

hey sorry i fell asleep! so the function is 1+x grams/cm^2 and this is what the pictures looks like that im given....